A reaction has an activation energy, Ea = 41.6 kJ mol-1 at 25 °C. By how many degrees must the temperature be raised in order to increase the reaction rate by a factor of 30.

  • A.

    2.0

  • B.

    6.0

  • C. 29

  • D.

    374

  • E.

    76

The following reaction takes place in an acidic solution

As4O6(s) + MnO4-(aq) → AsO43-(aq) + Mn2+(aq)

In the simplest balanced equation, the coefficients for each listed species are

As4O6 MnO4- H+

  • A.

    1,3,12

  • B. 3,5,36

  • C.

    5,8,36

  • D.

    5,8,30

  • E.

    2,3,5
Answer #1

1)
Given:
T1 = 25 oC
=(25+273)K
= 298 K
K2/K1 = 30/1
Ea = 41.6 KJ/mol
= 41600 J/mol

use:
ln(K2/K1) = (Ea/R)*(1/T1 - 1/T2)
ln(30/1) = (41600.0/8.314)*(1/298 - 1/T2)
3.4012 = 5003.6084*(1/298 - 1/T2)
T2 = 374 K
= (374-273) oC
= 101 oC

Increase in temperature = T2 - T1
= 101 oC - 25 oC
= 76 oC
Answer: E

Only 1 question at a time please